# Are LLMs effective psychological assessors? Leveraging adaptive RAG for interpretable mental health screening through psychometric practice

### v1

@@ Appendix B Dense Retrieval Models
Appendix C Benchmark Models
@@
Appendix D Ablation Study
D.1 The Impact of the Adaptive $k^{*}$ Dense Retrieval Approach
@@
D.2 The Impact of Different Prompting Strategies
In Table 8, we evaluate whether the use of direct prompting is statistically better than CoT across different metrics w.r.t. eRisk 2019 dataset collection. We perform both parametric, t-test, and non-parametric, Mann-Whitney U test (in both cases, we test whether one population mean is statistically greater than the other, $\alpha=0.05$). Open-source models perform well only in direct prompt contexts, while they perform poorly when the prompting technique is CoT, tending to overestimate questionnaire scores. Regarding closed-source LLMs, we observe that there isn’t always a significant difference between the two prompting approaches. For most metrics, with respect to gpt-4o-mini, the difference is significant only when using t-test.
D.3 The Impact of Different Retrieval Approaches
We also observe that some dense retrieval models perform better globally compared to others (see Figure 3), while some perform better only with respect to a subset of LLMs. In Figure 5, we examine the distribution of LLMs scores across different retrieval approaches and their rankings for each metric w.r.t. eRisk 2019 dataset collection. While no single approach consistently outperforms the others across all metrics, we can observe that some retrievers (msmarco-MiniLM-L-6-v3, msmarco-MiniLM-L-12-v3, distillbert-v4, and bge-large) generally achieve better rankings based on median values.
Figure 5: Distribution of adaptive RAG scores conditioned on different retrieval approaches and their rankings for each metric on the eRisk 2019 dataset. Rankings are based on median scores, with mean scores used as a tiebreaker in case of equal medians.

### v2

@@ Appendix C Dense Retrieval Models
Appendix D Benchmark Models
@@
Appendix E Ablation Study
E.1 Justification of the RAG approach
To further justify our retrieval-based approach, aka aRAG, we compare the performance of the most effective DCHR configurations using two closed-source LLMs against their performance when directly prompted to answer BDI-II items without filtering relevant posts. Specifically, we compare Claude 3.5 Sonnet + Direct prompt combined with T5 and gpt-4o-mini + Direct prompt combined with MiniLM-L12-v3. Given that some users have a significantly high number of posts, we input all posts that fit within each LLM’s context window based on timestamp order. We focus on the questionnaire-level metrics, DCHR and ADODL, which allow us to assess the accuracy of both approaches in determining depression severity. As shown in Figure 6, the aRAG approach consistently outperforms its no-retrieval counterpart across both collections and LLMs. Notably, we observe particularly wide performance gaps when using gpt-4o-mini on the eRisk 2019 collection (55.0% vs 35.0% for DCHR and 85.5% vs 82.0% for ADODL) and Claude on the 2020 edition (47.1% vs 31.4% for DCHR and 83.9% vs 80.9% for ADODL).
Figure 6: Performance comparison between aRAG and no-retrieval approaches for closed-source LLMs across DCHR and ADODL metrics on eRisk 2019 and 2020 datasets
E.2 The Impact of the Adaptive $k^{*}$ Dense Retrieval Approach
@@
E.3 The Impact of Different Prompting Strategies
In Table 7, we evaluate whether the use of direct prompting is statistically better than CoT across different metrics w.r.t. eRisk 2019 dataset collection. We perform both parametric, t-test, and non-parametric, Mann-Whitney U test (in both cases, we test whether one population mean is statistically greater than the other, $\alpha=0.05$). Lightweight open-source models (Phi-3-mini and Phi-3.5-mini) perform well only in direct prompt contexts, while they perform poorly when the prompting technique is CoT, tending to overestimate questionnaire scores (difference is statistically significant for ADODL, AHR, and ACR metrics).
For closed-source (gpt-4o-mini, Claude-3.5-Sonnet) and large-scale open-source LLMs (Qwen 2.5 70B, DeepSeek V3), the performance difference between prompting strategies varies. Specifically, gpt-4o-mini shows significant differences only under t-test across most metrics. Claude-3.5-Sonnet and DeepSeek V3 demonstrate significant differences solely in ADODL according to t-test, while Qwen 2.5 70B shows no statistically significant differences across any metric.
E.4 The Impact of Different Retrieval Approaches
We also observe that some dense retrieval models perform better globally compared to others (see Figure 3), while some perform better only with respect to a subset of LLMs. In Figure 7, we examine the distribution of LLMs scores across different retrieval approaches and their rankings for each metric w.r.t. eRisk 2019 dataset collection (after removing anomalous values given by lightweight open-source models in combination with CoT prompting strategies). While no single retrieval approach consistently outperforms the others across all metrics, we can observe that some retrievers (msmarco-MiniLM-L-6-v3, msmarco-MiniLM-L-12-v3, distillbert-v4, and bge-large) generally achieve better rankings on average.
Figure 7: Distribution of adaptive RAG scores conditioned on different retrieval approaches and their rankings for each metric on the eRisk 2019 dataset. Rankings are based on mean scores (shown as diamonds). The color gradient from brown to teal indicates performance ranking, with darker teal representing better performance.